OpenStudy (blaque678):

w^2+16w+?....I need help find the perfect square

OpenStudy (mathmate):

Expand (w+8)^2 and you will get the answer.

Parth (parthkohli):

\[\textbf{HINT}\]\[a^2 + 2ab + b^2 = (a + b)^2\]So,\[2ab =16w \iff ab = 8w \iff a = 8, b=w.\]

Parth (parthkohli):

The opposite...\[a = w,b=8\]

Parth (parthkohli):

So the last term is \(b^2 \) a.k.a \(8^2\)